### v1

# Hypergraph Neural Network with State Space Models for Node Classification

Abstract
In recent years, graph neural networks (GNNs) have gained significant attention for node classification tasks on graph-structured data. However, traditional GNNs primarily focus on adjacency relationships between nodes, often overlooking the rich role-based characteristics that are crucial for learning more expressive node representations. Existing methods for capturing role-based features are largely unsupervised and fail to achieve optimal performance in downstream tasks. To address these limitations, we propose a novel hypergraph neural network with state space model (HGMN), that effectively integrates role-aware representations into GNNs and state space model. HGMN utilizes hypergraph construction techniques to model higher-order relationships and combines role-based and adjacency-based representations through a learnable mamba transformer mechanism. By leveraging two distinct hypergraph construction methods—based on node degree and neighborhood levels—it strengthens the connections among nodes with similar roles, enhancing the model’s representational power. Additionally, the inclusion of hypergraph convolution layers enables the model to capture complex dependencies within hypergraph structures. To mitigate the over-smoothing problem inherent in deep GNNs, we incorporate a residual network, ensuring improved stability and better feature propagation across layers. Extensive experiments conducted on one newly introduced dataset and four benchmark datasets demonstrate the superiority of HGMN. The model achieves significant performance improvements on node classification tasks compared to state-of-the-art GNN methods. These results highlight HGMN’s ability to provide enriched node representations by effectively embedding role-based features alongside adjacency information, making it a versatile and powerful tool for a variety of graph-based learning applications.
keywords:
Graph neural networks, Hypergraph, State space model, Node classification.
journal: Elsevier
[…]
[inst1]organization=Department of Mathematics, Indian Institute of Technology Indore,addressline=Simrol, city=Indore, postcode=453552, state=Madhya Pradesh, country=India
1 Introduction
Graph-based modeling has become a prevalent approach for analyzing complex data structures and relationships, including applications like social networks [1], molecular interactions [2], and brain connectivity [3]. Graph neural networks (GNNs) have demonstrated superior performance over traditional neural networks in a variety of applications [4]. Prominent examples of GNNs include models such as GCN [5], GAT [6], and GraphSAGE [7]. These networks are designed to capture the spatial characteristics inherent in graph data, enhancing both node-level and graph-level representations. A key principle underlying most GNNs is the aggregation of information from neighboring nodes, a concept often referred to as message passing [8]. This approach focuses on leveraging features from adjacent or connected nodes within the graph to improve representation learning.
In a graph, nodes typically represent different social roles or functions, and nodes that share similar roles or functions are not always directly connected. Traditional GNNs, such as GCN [5], GAT [6], and GraphSAGE [7], generate feature embeddings for central nodes by aggregating information from their neighbors. While this approach captures local subgraph structures, it has notable limitations: (1) It faces challenges in generating node representations that capture detailed and role-specific features effectively, and (2) It encounters difficulties in linking nodes with similar roles that are distant or lack direct connections. A node’s role can reflect its real-world semantic function such as being a student, advisor, or manager, or its structural role within a network, such as an opinion leader or structural hole spanner [9]. Although unsupervised methods have been proposed for role representation learning [10, 11], their performance in downstream tasks remains suboptimal [7]. Furthermore, adjacency-based multi-layer GNNs frequently face over-smoothing problems, where nodes become indistinguishable after several layers of message passing.
Traditional GNNs primarily work with standard graphs that represent node adjacency while overlooking the role-based relationships between nodes. In contrast, hypergraphs [12] can be derived from conventional graph structures in multiple ways. Hypergraph adjacency encodes richer relationships among nodes and clusters, enhancing the effectiveness of aggregation processes in graph neural networks (NNs) for capturing both role-based and adjacency-based representations.
Recently, Graph Transformers have become a powerful tool due to their ability to effectively capture long-range interactions between nodes [13, 14]. Unlike traditional local message-passing methods that focus on edge-level neighborhood structures [5, 15], the attention mechanisms in Graph Transformers allow each node to interact with every other node in the graph, providing a more global perspective [16]. To enhance the development of Graph Transformers, the GraphGPS framework integrates structural/positional encodings (SE/PE), a message-passing NN (MPNN), and an attention module. Collectively, these components refine node and edge embeddings to facilitate downstream tasks. This modular design allows for the seamless incorporation of various attention mechanisms, offering significant flexibility for users [17].
The quadratic computing cost of the attention mechanism limits the usefulness of transformers for handling long sequences, despite their significantly enhanced computational capability. Furthermore, it might not always be optimal for long-range dependencies to fully capture attention or encode all contextual information. Empirical studies suggest that increasing context length often fails to improve the performance of many sequence models [18]. State space models (SSMs) have surfaced as an effective alternative to attention-based frameworks such as Transformers [19]. However, their performance falls short of Transformers due to constraints in capturing context compression that depends on the input, a drawback of their time-invariant transition mechanisms. Mamba has shown remarkable success in language modeling, exceeding the performance of Transformers of the same size and achieving results comparable to Transformers that are twice as large. These achievements have inspired recent efforts to adapt Mamba’s architecture for various data modalities [20, 21].
Building on the concept of hypergraphs [12], we introduce the hypergraph NN with state space model (HGMN), which integrates role-based representations into GNNs. HGMN leverages hypergraph construction methods to integrate node representations of roles and adjacency from higher-order viewpoints. We incorporate two methods for constructing hypergraphs that aggregate node information based on degree and neighborhood levels, reinforcing connections between nodes with similar roles within the hypergraph. HGMN integrates a mamba transformer mechanism with adaptive weights to merge adjacency-based and role-based representations. Through hypergraph convolution layers, the model learns hypergraph structures during training. A residual network is used to address the over-smoothing problem between convolutional layers. Compared to standard GNNs, HGMN enriches node representations by incorporating role characteristics and adopts a more effective aggregation approach. It is versatile and well-suited for node representation tasks involving both adjacency structures and role information.
The paper’s key highlights are as follows:
  1. 1.
    We propose a novel hypergraph NN with state space model (HGMN), that integrates role-based representations into graph NNs. HGMN leverages hypergraph construction techniques to effectively combine adjacency and role-based node representations from higher-order perspectives.
  2. 2.
    Two unique hypergraph construction methods are incorporated, based on node degree and neighborhood levels. These methods enhance role representation learning by ensuring stronger connections between nodes with similar roles in the hypergraph structure.
  3. 3.
    HGMN incorporates a learnable mamba transformer mechanism to efficiently merge role-based and adjacency-based node representations, enabling adaptive and robust representation learning.
  4. 4.
    By employing hypergraph convolution layers, HGMN learns hypergraph structures during training. To address the over-smoothing problem common in graph NNs, a residual network is incorporated, improving the stability and expressiveness of the model.
  5. 5.
    Compared to traditional GNNs, HGMN provides enriched node representations by embedding role characteristics and utilizing a more effective aggregation strategy. This makes the model versatile and applicable to a wide range of node representation tasks involving adjacency structures and role information.
  6. 6.
    HGMN effectiveness is validated through comprehensive experiments on four publicly available datasets. The results demonstrate that HGMN achieves significant performance gains, with absolute improvements of up to $12.1\%$ in node classification accuracy compared to existing GNN methods.
2 Preliminaries and related work
An undirected graph can be represented as $G=(V,E,A)$, where $V=\{v_{1},v_{2},\dots,v_{N}\}$ is the set of nodes, $A\in\mathbb{R}^{N\times N}$ is the adjacency matrix, and $E$ represents the set of edges. The role representation is denoted as $X_{r}\in\mathbb{R}^{N\times F_{r}}$, while $X_{a}\in\mathbb{R}^{N\times F_{a}}$ corresponds to the adjacency-based representation. Node labels are indicated by $Y\in\mathbb{R}^{N}$, where the nodes belong to $M$ distinct categories. Graph neural networks (GNNs) generate the node representations $Z\in\mathbb{R}^{N\times F_{h}}$ and use them to predict the labels of the nodes.
2.1 Graph Neural Network
Graph Neural Networks (GNNs), introduced in 2009, are designed to analyze data that is structured as graphs [22]. Over recent years, research interest in GNNs has grown significantly [23, 24]. This research can generally be divided into two main categories. The first focuses on applying GNNs to a variety of tasks, including knowledge representation [25], text generation [26], traffic prediction [27], and weather forecasting [28].
ChebNet [29] introduced a spectral approach using Chebyshev polynomials to extract features from graph-structured data. Building on this, GCN [5] simplified the method with feature propagation to aggregate adjacency information, improving computational efficiency. Shi et al. [30] introduced UniMP, a consolidated message-passing framework that combines feature and label propagation during both the training and inference stages. SAT [14] enhances the performance of existing GNNs by extracting subgraph representations, improving their overall effectiveness. These GNNs primarily rely on adjacency-based message-passing for feature aggregation.
Beyond adjacency-driven feature aggregation, some GNNs incorporate additional characteristics such as node identities or positional information, which are absent in traditional methods. For instance, ID-GNN [31] adopts an inductive approach to incorporate node identities into the message-passing process. Similarly, the position-aware GNN (PGNN) [32] focuses on generating embeddings that account for node locations. Li et al. [33] introduced GLOGNN and its enhanced version, GLOGNN++, which utilizes a coefficient matrix to enable neighborhood aggregation and capture inter-node correlations.
Xie et al. [34] developed the scale-aware GNN (SAGNN), which models cross-scale interactions for tasks such as few-shot semantic segmentation. At the same time, Xu et al. [35] introduced the graph wavelet NN (GWNN), which utilizes graph wavelet transforms to overcome the limitations of earlier spectral graph convolutional networks that depend on graph Fourier transforms. Although these approaches demonstrate strong performance, they do not directly tackle the connection between role-based information and higher-order feature aggregation within graph structures.
2.2 Hypergraph Learning
Hypergraph learning was first proposed as a propagation method on hypergraph structures. Thanks to its degree-free hyperedges, hypergraph learning enables the modeling of high-order data correlations beyond simple pairwise relationships [36]. Early works, such as those by Feng et al. [12] and Yadati et al. [37], were pioneers in integrating hypergraph structures into GNNs, marking a significant step in the evolution of GNNs. Bai et al. [38] advanced this direction by introducing hypergraph attention networks and hypergraph convolutional NNs to enhance representation learning capabilities. Zhang et al. [39] proposed the dual-channel hypergraph convolutional framework (DHCF) for collaborative filtering in recommendation systems, demonstrating hypergraphs’ utility in practical applications. Chien et al. [40] contributed a generalized framework for hypergraph NNs that dynamically chooses propagation methods suited to the characteristics of individual datasets. Yang et al. [41] further refined this concept by introducing an attention-based hypergraph NN that prioritizes essential features during learning. Despite these advancements and the broad application of hypergraph NNs across various domains, their potential for handling node representations rich in role-based features remains largely unexplored.
2.3 State Space Models
State space models (SSMs), which range from hidden Markov models to recurrent NNs (RNNs) in contemporary deep learning, use recurrent updates over sequences through hidden states. Using a recurrent scanning approach, SSMs maintain context within hidden states and update outputs by integrating these states with incoming inputs. structured state space models (S4) improve computational efficiency by leveraging reparameterization techniques, providing a streamlined alternative to traditional attention mechanisms [42].
Recent developments in linear-time attention models based on S4 include innovations such as hungry hungry hippos (H3) [43], hyena [44], and receptance weighted key value RWKV [45]. Building on this foundation, Mamba integrates a data-driven selection mechanism within S4, allowing for efficient capture of long-range dependencies as the sequence lengths increase [18]. Mamba achieves linear-time complexity in long-sequence modeling and outperforms Transformers across various benchmarks. Moreover, its adaptability extends beyond sequential data, demonstrating effectiveness in tasks like image segmentation by enhancing long-range contextual understanding [46].
Figure 1: Overview of the proposed framework
Figure 2: Simple graph vs. Hypergraph. (a) A simple graph along with its vertex-edge incidence matrix. (b) A hypergraph and its corresponding vertex-edge incidence matrix.
3 Hypergraph Neural Network with State Space Model (HGMN)
Figure 1 depicts the structure of the HGMN framework, comprising three primary components. Initially, the graph structure is encoded to derive the role and adjacency representations using Node2vec [12] and Graphwave [11], respectively. A hypergraph is then constructed from the input graph. These representations are combined into a unified embedding through an attention mechanism. Finally, a multi-layer hypergraph convolutional network utilizes the fused embeddings and hypergraph adjacency matrix for iterative computations.
3.1 Construction of Hypergraph
This subsection outlines two detailed methods for constructing hypergraphs. The resulting hypergraphs are utilized in the following hypergraph convolutional networks to perform feature aggregation.
3.1.1 Description of the Hypergraph
In a standard graph, edges connect pairs of nodes, while in a hypergraph, a hyperedge can connect several nodes that share common characteristics. This allows a simple graph to be restructured into a hypergraph. A hypergraph $G_{H}=(V_{H},E_{H},W_{H})$ is defined with $H$ represented as $\theta(v,e)$, where $\theta(v,e)$ is either 0 or $1$, signifying whether node $v$ belongs to a hyperedge $e$. $V_{H}$ denotes the set of nodes, $E_{H}$ represents the set of hyperedges, and $W_{H}$ is the hyperedge weight matrix. The degree of a hyperedge is given by:
$d(e_{i})=\sum_{v_{i}\in V}\theta(v_{i},e_{i}).$ (1)
The node degree is computed as:
$d(v_{i})=\sum_{e_{i}\in E}\theta(v_{i},e_{i})w(e_{i}),$ (2)
where the degree matrices for nodes and hyperedges are given by $D_{v}$ and $D_{e}$, respectively. The techniques for hypergraph construction in HGMN are shown in Fig.2.
3.1.2 Node-Link Hypergraph
This approach creates a hypergraph $G^{l}_{h}$ by leveraging the neighboring connections within a graph $G$. In $G^{l}_{h}$, the neighboring nodes of a central node are clustered together to create a hyperedge. $H^{l}\in\mathbb{R}^{N\times N^{l}_{E}}$ represents the resulting hypergraph structure, where $N$ denotes the total number of nodes in $G$, and $N^{l}_{E}$ represents the number of hyperedges. The degree of nodes and hyperedges within $H^{l}$ are denoted by $D^{l}_{v}$ and $D^{l}_{e}$, respectively.
3.1.3 Hypergraph Based on Node Degree
According to their degrees, nodes are grouped to form a hypergraph $G^{d}_{h}$. Nodes sharing the same degree are combined into a single hyperedge. The resulting hypergraph structure is represented as $H^{d}\in\mathbb{R}^{N\times N^{d}_{E}}$, where $N$ is the total number of nodes, and $N^{d}_{E}$ corresponds to the number of unique node degrees. The degree matrices for nodes and hyperedges in $H^{d}$ are denoted as $D^{d}_{v}$ and $D^{d}_{e}$, respectively.
3.2 Mamba workflow
HGMN incorporates an attention mechanism to enable the adaptive integration of role-based and adjacency-based representations, enhancing performance in downstream tasks. Role features are embedded into $\hat{X}_{r}$ to achieve this. Additionally, HGMN leverages a structured state space model (SSM) attention mechanism for effective sequence modeling, ensuring the fusion of dynamic role and adjacency information.
SSM [18], a sequential model, employs a linear ordinary differential equation (ODE) to transform an input sequence $x(t)\in\mathbb{R}^{N}$ into an output sequence $y(t)\in\mathbb{R}^{N}$ through a latent state $h(t)\in\mathbb{R}^{N}$:
$h^{\prime}(t)=Ah(t)+Bx(t),$
$y(t)=Ch(t),$ (3)
where the state matrix is represented by $A\in\mathbb{R}^{N\times N}$, the input matrix, and the output matrix by $B$ and $C\in\mathbb{R}^{N}$. To determine $y(t)$ at time $t$, the latent state $h(t)$ must be solved, which is analytically challenging. Considering that real-world data is often discrete, we discretize the ODE as follows:
$h_{t}=\bar{A}h_{t-1}+\bar{B}x_{t},\quad y_{t}=Ch_{t},$ (4)
where the discretized state parameters are $\bar{A}=\exp(\Delta\cdot A)$ and $\bar{B}=(\Delta\cdot A)^{-1}(\exp(\Delta\cdot A)-I)\cdot\Delta B$, with $\Delta$ standing for the step size.
In HGMN, the SSM attention mechanism utilizes these state transitions to integrate and propagate role and adjacency-based features effectively. By calculating the output sequence $y_{t}$, the model learns dynamic interactions between roles and structural features, enhancing its ability to represent higher-order relationships within the hypergraph. The attention coefficient $y_{r}$ is computed as follows:
$X_{f}=\hat{y}_{r}\hat{X}_{r}+\hat{y}_{a}\hat{X}_{a}.$ (5)
Eq (5) is used to generate the resulting fused embedding, $X_{f}\in\mathbb{R}^{N\times F_{h}}$, which is used as the input feature set for the hypergraph convolutional networks.
3.3 Hypergraph Convolution Layer
The fused features obtained from the attention mechanism (eq (5)) are the initial input of the hypergraph convolutional network. The node fusion features $X$, processed through a hypergraph convolutional network (HGMN), are computed as follows:
$X^{(l)}=\sigma\left(D_{v}^{-1}HW_{H}D_{e}^{-1}H^{T}D_{v}^{-1}X^{(l-1)}\delta^{(l-1)}\right)$ (6)
Here:
  • 1.
    $W_{H}$ represents the weight matrix associated with the hyperedges.
  • 2.
    $D_{e}$ and $D_{v}$ represent the degree matrices for hyperedges and nodes, respectively.
  • 3.
    $H$ denotes the matrix representing the hypergraph structure.
  • 4.
    $X^{(l-1)}$ represents the input at the $l$-th layer, and $X^{(l)}$ corresponds to its output.
  • 5.
    $\delta^{(l-1)}$ denotes the learnable parameters for the $l$-th layer.
  • 6.
    $\sigma$ is the nonlinear activation function applied after the convolution operation.
3.4 Residual Network
A residual connection is incorporated to mitigate the over-smoothing problem often encountered in GNNs. This approach involves down-sampling the raw input features and combining them with the output from the last HGMN layer. The process is defined as:
$J=X_{1}W_{\text{res}}+X_{\text{last}},$ (7)
where $W_{\text{res}}\in\mathbb{R}^{F_{h}\times F_{h}}$ is the weight matrix for transformation, $X_{1}\in\mathbb{R}^{N\times F_{h}}$ represents the raw input features, $X_{\text{last}}\in\mathbb{R}^{N\times F_{h}}$ is the output from the last HGMN layer, $J\in\mathbb{R}^{N\times F_{h}}$ is the resulting representation after applying the residual connection.
3.5 Classification
The representation $J$ is passed through a fully connected layer followed by a softmax function to generate predictions $\hat{Y}$:
$\hat{Y}=\text{softmax}(b_{m}+JW_{m}),$ (8)
where $b_{m}\in\mathbb{R}^{M}$ represents the bias term and $W_{m}\in\mathbb{R}^{F_{h}\times M}$ is the weight matrix.
The model’s loss function is defined as:
$L=\lambda\|\delta\|^{2}-\sum_{v\in V}\sum_{i=1}^{M}Y\ln\hat{Y}$, (9)
where $\lambda$ is the regularization parameter, which helps prevent overfitting, $\delta$ denotes the model parameters.
The algorithm of the proposed HGMN is briefly described in Algorithm 1.
Algorithm 1 HGMN
Input: Let $G=(V,E,A)$ represent the graph.
Output: Return node embeddings $J$.
1: Set the model parameters $\theta$ to random initial values.
2: Generate the hypergraph $G^{\mathcal{L}}$ or $G^{\mathcal{D}}$ using the adjacency matrix AA as the foundation.
3: Set up the role-specific embeddings $X_{r}$ and the adjacency-driven embeddings $X_{a}$.
4:while the model has not converged do
5:  Compute the fusion embeddings using the mamba transformer mechanism.
6:  Utilize multiple hypergraph convolution layers to compute $X_{\text{last}}$.
7:  Use the residual network to derive the final embeddings $J$.
8:  Update model parameters $\theta$ by minimizing the cross-entropy loss.
9:end while
10: Output the final node representations $J$.
4 Experiments
This section presents a series of comprehensive experiments across datasets of varying scales and methods to evaluate the effectiveness of the proposed HGMN model.
Datasets: The model’s performance is evaluated on six publicly available datasets. Detailed descriptions are as follows:
  1. 1.
    ENZYMES: This dataset, derived from the BRENDA enzyme database [47], consists of $600$ tertiary protein structures. We focus on enzyme graphs with more than $90$ nodes, specifically those indexed as $118$, $123$, $295$, $296$, and $297$. Each graph consists of two categories of nodes. For brevity, ENZYMES118 is denoted as E118, with similar abbreviations applied to the other subsets.
  2. 2.
    Internet Industry Partnerships (IIP) [48]: This dataset represents partnerships in the Internet industry. Nodes correspond to companies, and edges denote announced collaborations such as joint ventures or strategic alliances. The nodes are categorized into three tags: Content, Infrastructure, and Commerce.
  3. 3.
    TerroristRel: Obtained from the PIT repository [49], this dataset provides details about terrorists and their connections. Each relationship is characterized by a binary vector.
  4. 4.
    Cora [50]: This dataset contains $2708$ scientific publications related to machine learning, with nodes categorized into seven distinct groups.
  5. 5.
    Citeseer [50]: Comprising $3327$ scientific papers, nodes in this dataset are classified into six groups.
  6. 6.
    Pubmed [50]: Comprises $19,717$ scientific articles on diabetes from the Pubmed database, with nodes divided into three categories.
In the ENZYMES dataset, nodes are categorized into primary and secondary types as outlined in [47]. The IIP dataset highlights extensive market connections between established corporations and emerging startups [48]. The TerroristRel dataset reflects disparities in the influence of different factions [49]. Meanwhile, in Cora, Citeseer, and Pubmed, key research papers serve as central references, being frequently cited by others [50].
Baselines: The performance of HGMN is benchmarked against the following well-established methods:
  1. 1.
    Role: Node role representations generated using GraphWave [11] are utilized as input features for classification with an SVM employing a Radial Basis Function (RBF) kernel [51].
  2. 2.
    Adj: Node adjacency features derived from Node2vec [52] serve as input to an SVM classifier.
  3. 3.
    Role+Adj: A combination of role and adjacency features is fed into an SVM classifier for evaluation.
  4. 4.
    Deepwalk: Deepwalk [53] employs random walks and the Skip-Gram model [54] to learn node embeddings for graph data, which are then classified using the same SVM.
  5. 5.
    Role+GCN: Role-based embeddings obtained via GraphWave [11] are used as input features for a Graph Convolutional Network (GCN) to perform classification.
  6. 6.
    Struc2vec: Struc2vec [55] produces embeddings by leveraging vertex similarity obtained through role-biased Markov walks.
  7. 7.
    Adj+GCN: Adjacency embeddings from Node2vec [52] are applied as input to GCN for node classification tasks.
  8. 8.
    Role+GAT: Role embeddings from GraphWave [11] are utilized as input features for the Graph Attention Network (GAT) in the classification process.
  9. 9.
    Role+HGCN: Role features extracted by GraphWave [11] are fed into a hypergraph convolutional network (HGCN) [12] for classification.
  10. 10.
    Adj+GAT: Node2vec [52] adjacency embeddings serve as input features for GAT-based classification.
  11. 11.
    Adj+HGCN: For classification, the same HGCN model employs adjacency embeddings generated by Node2vec [52] as input.
  12. 12.
    StrucGCN [56]: StrucGCN model leverages a structural matrix based on topological similarity to enhance structural information learning, enabling effective non-local aggregation and improved performance on heterophily graphs.
  13. 13.
    AGNN [57]: AGNN integrates graph convolutional layers (GCL) with Graph Embedding Layers (GEL) derived from Laplacian-regularized optimization to mitigate over-smoothing by alternating between low- and high-order feature spaces.
Parameter Settings: In HGMN, we utilize two hypergraph convolutional layers followed by a fully connected layer. In Node2vec, the parameters $p$ and $q$ are both set to $1.0$. Adam optimizer is used with an initial learning rate of $0.003$, and the weight decay coefficient is set to 5e45e-4. For the baseline methods, the parameters are configured using their default settings.
Table 1: Performance of the proposed HGMN models in node classification, compared to the baseline models.
Model $\downarrow$ Dataset $\rightarrow$ E118 E123 E295 E296 E297 IIP TerroristRel Cora Citeseer Pubmed
Deepwalk $54.76\pm 8.38$ $46.35\pm 12.44$ $45.73\pm 12.69$ $42.66\pm 9.05$ $45.70\pm 11.99$ 60.85 51.77 30.09 33.60 33.20
Struc2vec $55.38\pm 10.02$ $46.01\pm 13.11$ $45.16\pm 12.81$ $42.65\pm 9.47$ $46.35\pm 10.60$ 61.24 51.91 30.06 - -
Role $54.55\pm 9.89$ $46.23\pm 13.15$ $45.21\pm 11.80$ $2.87\pm 9.46 44$$4.41\pm 12.28$ 60.62 52.11 30.23 19.62 39.05
Adj $53.66\pm 9.99$ $7.23\pm 13.37 44$$4.86\pm 12.39$ $42.02\pm 10.44$ $47.31\pm 12.39$ 60.79 51.91 30.02 19.40 71.65
Adj+GCN $58.72\pm 8.33$ $61.23\pm 14.07$ $61.67\pm 12.73$ $59.32\pm 12.94$ $64.04\pm 12.31$ 58.73 55.92 71.74 51.84 70.72
Role+GAT $54.72\pm 9.58$ $52.53\pm 14.79$ $58.55\pm 15.59$ $54.94\pm 13.78$ $58.23\pm 13.80$ 60.18 53.24 40.51 41.24 35.70
Role+Adj $52.07\pm 8.44$ $64.24\pm 13.07$ $45.91\pm 13.50$ $41.15\pm 9.67$ $45.37\pm 11.10$ 59.50 51.33 30.02 19.35 39.77
Adj+GAT $56.55\pm 9.86$ $56.89\pm 14.17$ $59.31\pm 14.91$ $56.23\pm 12.59$ $60.09\pm 13.84$ 60.08 53.25 71.96 51.82 71.90
Role+GCN $58.55\pm 9.44$ $61.74\pm 14.11$ $61.93\pm 12.89$ $59.74\pm 13.10$ $59.56\pm 12.26$ 60.23 55.17 38.74 39.38 38.53
Adj+HGCN $(\mathcal{D})$ $55.28\pm 9.58$ $54.51\pm 14.53$ $51.55\pm 13.34$ $51.05\pm 13.68$ $51.55\pm 14.39$ 58.95 53.78 30.06 24.89 35.20
Role+HGCN $(\mathcal{L})$ $57.10\pm 9.88$ $54.10\pm 15.51$ $58.14\pm 14.52$ $56.88\pm 12.82$ $58.24\pm 12.80$ 60.18 53.73 44.16 42.39 41.21
Adj+HGCN $(\mathcal{L})$ $58.07\pm 8.63$ $57.10\pm 14.61$ $58.48\pm 13.79$ $57.97\pm 13.13$ $58.94\pm 13.24$ 58.64 53.96 72.26 51.74 71.17
Role+HGCN $(\mathcal{D})$ $54.31\pm 9.63$ $52.68\pm 14.59$ $50.89\pm 13.47$ $49.20\pm 14.10$ $50.89\pm 14.14$ 60.41 53.42 38.74 24.70 28.54
AGNN $54.55\pm 3.69$ $63.23\pm 5.27$ $59.87\pm 2.86$ $58.76\pm 6.25$ $60.53\pm 4.25$ $61.42$ $53.23$ $70.52$ $51.87$ $71.87$
StrucGCN $54.62\pm 7.48$ $63.89\pm 10.70$ $60.38\pm 4.79$ $59.35\pm 11.26$ $62.48\pm 7.82$ $61.08$ $52.48$ $71.89$ $51.32$ $71.50$
HGMN $(\mathcal{D})$ $62.89\pm 5.72$ $65.42\pm 4.42$ $62.47\pm 6.42$ $63.87\pm 8.64$ $65.00\pm 8.42$ 62.62 59.72 71.56 52.69 72.89
HGMN $(\mathcal{L})$ $\mathbf{63.23}\pm 4.82$ $\mathbf{67.59}\pm 4.28$ $\mathbf{63.20}\pm 6.72$ $\mathbf{65.42}\pm 5.94$ $\mathbf{67.25}\pm 8.59$ $61.77$ $\mathbf{60.42}$ $\mathbf{73.14}$ $50.78$ $\mathbf{72.99}$
AI 4.51 3.35 1.27 5.68 3.21 1.38 4.50 0.88 0.85 1.09
IR 7.68% 5.21% 2.05% 9.50% 5.01% 2.25% 8.04% 1.22% 1.64% 1.52%
represents the proposed models.
AI represents the absolute improvement. IR represents the improvement ratio.
4.1 Classification of Node
The node classification task is employed to evaluate the performance of HGMN and the baseline methods. HGMN includes two alternative models: HGMN($\mathcal{D}$), which is based on node degree, and HGMN($\mathcal{L}$), which relies on node-link characteristics. Both the TerroristRel and ENZYMES datasets have imbalanced class distributions. An equal distribution of these datasets may cause the model to prioritize classes with more abundant samples. To address this, we performed random sampling for the training and testing sets, ensuring that the ratio of larger to smaller node sets remained within $1:0.33$ for imbalanced datasets. For Cora, Citeseer, and Pubmed, only structural graph features were used for classification, excluding the original features. As a result, the performance of some baselines may differ from the results reported in [5]. Due to the smaller size of the Adv-Stu and ENZYMES graph structures, the prediction accuracy tends to vary more. To ensure reliable results, we conducted multiple runs: $100$ trials for the Adv-Stu and ENZYMES datasets, $50$ trials for TerroristRel and IIP, and $10$ trials for Citeseer, Cora, and Pubmed, all with consistent data preprocessing. Since the graph structures of IIP, TerroristRel, Citeseer, Pubmed, and Cora are larger, we expect optimal performance on these datasets. We present both the average and best micro-F1 scores for evaluation.
Table 2: Maximum Performance of the proposed HGMN models in node classification, compared to the baseline models.
Model $\downarrow$ Dataset $\rightarrow$ E118 E123 E295 E296 E297 IIP TerroristRel Cora Citeseer Pubmed
Deepwalk 65.23 75.78 80.98 75.19 82.72 72.73 71.11 33.6 23.4 33.8
Struc2vec 66.45 76.35 75.38 76.27 83.02 74.24 70.59 33 - -
Role 62.27 79.82 77.98 73.38 84.98 72.73 67.57 32.6 23.1 39.1
Adj 70.48 82.84 80 78.01 87.9 72.73 67.57 32.9 22.9 72.5
Adj+GCN 72.58 80.89 78.55 79.74 82.98 70.45 69.39 73.5 52.7 70.9
Role+GAT 68.23 76.45 77.44 75.97 83.11 72.73 65.85 42.8 42.9 41.8
Role+Adj 67.92 76.35 81.11 70.32 83.85 72.73 70 32.1 22.8 43.8
Adj+GAT 69.49 81.78 78.81 75.11 86.62 72.73 65.85 73.4 53.3 72.4
Role+GCN 70.14 78.42 76 79.38 82.8 72.73 71.43 41.4 39.9 40.8
Adj+HGCN $(\mathcal{D})$ 70.82 77.87 74.59 78.92 88.25 70.45 65.85 32.9 25.8 37.1
Role+HGCN $(\mathcal{L})$ 72.63 79.62 78.26 82.52 86 72.73 65.85 48.1 43.4 42.1
Adj+HGCN $(\mathcal{L})$ 72.98 75 78.26 80.02 81.5 70.45 65.85 73.1 53.9 73.7
Role+HGCN $(\mathcal{D})$ 71.54 80.23 82.72 78.97 86.52 75 65.85 41.31 25.2 40.7
AGNN $70.89$ $82.56$ $80.76$ $80.48$ $85.29$ $73.52$ $71.42$ $69.59$ $52.82$ $70.49$
StrucGCN $71.46$ $82.39$ $82.46$ $81.36$ $87.43$ $73.97$ $70.78$ $72.87$ $53.46$ $72.19$
HGMN $(\mathcal{D})$ 72.41 83.33 83.33 81.82 90 68.18 64.29 63.4 49.3 71.2
HGMN $(\mathcal{L})$ 78.42 80.49 79.48 82.75 89.29 76.78 72.82 74.42 54.63 75.72
AI 5.44 0.49 0.61 0.23 2.10 2.54 1.39 0.92 0.73 2.02
IR 7.45% 0.59% 0.74% 0.28% 2.39% 3.42% 1.95% 1.25% 1.35% 2.74%
represents the proposed models.
AI represents the absolute improvement. IR represents the improvement ratio.
The results are presented in Tables 1 and 2, where Micro-F1 is used as the evaluation metric for the node classification task. Micro-F1 is widely used in classification tasks to assess model performance, as it provides a balanced measure of precision and recall, which reflects the overall effectiveness of the model in distinguishing between classes. It is calculated as:
$\text{Micro-F1}=\frac{2\times(\text{Recall}_{\text{micro}}\times\text{Precision}_{\text{micro}})}{\text{Recall}_{\text{micro}}+\text{Precision}_{\text{micro}}}$ (10)
To assess the improvement of HGMN over the baselines, we calculate the Absolute Improvement (AI) as P1P2P1-P2, where $P1$ denotes the top performance of HGMN, while $P2$ indicates the best performance among the existing methods. The Improvement Ratio is then computed as:
$\text{Improvement Ratio (IR)}=\frac{(P1-P2)}{P2}\times 100\%$ (11)
The results show that HGMN outperforms all the other baseline models across all datasets, achieving the highest performance in every case. In particular, HGMN demonstrates significant improvements over the baselines, with large margins in most cases. Incorporating neighbor links, which represent the local subgraph structure of each node, proves especially effective in datasets such as IIP, ENZYMES, Cora, TerroristRel, Pubmed, and Citeseer. These datasets benefit from the rich structural information provided by node connections, allowing HGMN to better capture the relationships and improve classification accuracy.
Figure 3: Effect of the learning rate $\lambda$ on the performance of the proposed HGMN model.
Table 3: Ablation of proposed HGMN model.
Model $\downarrow$ Dataset $\rightarrow$ E118 E123 E295 E296 E297 IIP TerroristRel Cora Citeseer Pubmed
HGMN $(\mathcal{D})$ 62.89 65.42 62.47 63.87 65.00 62.62 59.72 71.56 52.69 72.89
HGMN $(\mathcal{D})$/residual 59.38 64.48 64.04 63.09 63.89 59.86 60.07 32.27 25.42 43.90
HGMN $(\mathcal{D})$/mamba block 61.31 66.93 66.06 65.46 64.89 60.45 60.09 70.90 50.35 73.61
HGMN $(\mathcal{L})$ 63.23 67.59 63.20 65.42 67.25 61.77 60.42 73.14 50.78 72.99
HGMN $(\mathcal{L})$/residual 63.79 69.39 70.85 69.69 69.41 61.27 60.26 72.32 50.93 74.20
HGMN $(\mathcal{L})$/mamba block 63.28 69.20 70.39 68.07 68.57 60.91 58.62 71.89 51.63 74.78
The bold value indicates the highest performance achieved among the models.
4.2 Ablation Study
We performed a comprehensive ablation study to investigate the impact of each component in the HGMN model. Table 3 presents the summarized results. We tested modified versions of the model by systematically removing specific components, such as the residual network and the mamba block. These modifications are denoted as HGMN/residual and HGMN/mamba block, respectively. The analysis reveals that eliminating either the residual network or the mamba block significantly reduces the model’s accuracy across multiple datasets. This highlights the essential role of both components in improving the overall performance of HGMN. In particular, the residual network facilitates the integration of initial representations with global network information, enabling the model to achieve superior global optimization. This effect is especially pronounced in HGMN $\mathcal{D}$, where removing the residual network leads to a dramatic decline in performance on large citation network datasets such as Cora, Citeseer, and Pubmed.
In contrast, HGMN $\mathcal{L}$ demonstrates a relatively smaller dependency on the residual network. Instead, the mamba block plays a more prominent role in this variant, allowing the model to better capture the intricate relationships among nodes. This suggests that the mamba block is particularly effective in leveraging the local structural information in graph networks. Overall, the results confirm that HGMN outperforms its variants and other baseline models by significant margins across all tested datasets. The inclusion of both the residual network and mamba block ensures that HGMN achieves robust and consistent performance, particularly in large and complex network structures. These findings highlight the strong design and adaptability of HGMN in handling diverse graph-based classification tasks.
4.3 Sensitivity Analysis
The graph in Fig. 3 highlights how the learning rate ($\lambda$) influences the performance of the proposed HGMN model across different datasets. The sensitivity of the model to changes in the learning rate varies depending on the dataset. For ENZYMES and IIP, a relatively higher learning rate tends to yield better results, suggesting that these datasets benefit from faster convergence during optimization. The inherent characteristics of these datasets, such as their graph structure or node distribution, might make them more adaptable to quicker updates provided by larger learning rates.
In contrast, datasets such as Cora, Citeseer, and Pubmed demonstrate improved performance when a smaller learning rate is used. These datasets are typically larger and more complex, where a lower learning rate helps in achieving more stable and fine-tuned updates to the model parameters. This is particularly important in avoiding overshooting during training and ensuring the optimization process gradually converges to a better solution. The results emphasize the importance of carefully tuning the learning rate based on the dataset characteristics to maximize the model’s performance. Selecting an appropriate learning rate is critical to balancing the trade-off between training speed and achieving higher predictive accuracy.
5 Conclusion
In many graph datasets, nodes that share similar roles are not necessarily directly linked, which limits the effectiveness of traditional graph neural networks (GNNs) that focus primarily on adjacency relationships while neglecting role-based characteristics. To address this gap, we introduced HGMN, a hypergraph neural network with state space model, which integrates role-aware and adjacency-based representations through hypergraph construction techniques and a learnable mamba transformer mechanism. Comprehensive experiments across multiple datasets showed that HGMN delivers state-of-the-art performance, greatly surpassing baseline models. These results highlight the effectiveness of HGMN in bridging role and adjacency representations, offering a fresh perspective for designing advanced GNN architectures. Additionally, HGMN’s scalable design makes it adaptable as a broad framework for representing graphs in networks that have role characteristics. However, many real-world graphs are heterogeneous, with node roles and adjacency relationships shaped by more complex dependencies and constraints. Currently, HGMN is not optimized to address the complexities associated with heterogeneous graphs. Future work will focus on extending HGMN to accommodate heterogeneous graph structures and explore its applicability to a broader range of graph-based tasks.
Acknowledgement
This study receives support from the Science and Engineering Research Board (SERB) through the Mathematical Research Impact-Centric Support (MATRICS) scheme Grant No. MTR/2021/000787. The authors gratefully acknowledge the invaluable support provided by the Indian Institute of Technology Indore.